OpenStudy (anonymous):

Please someone can help me solve this question #11. Thanks

OpenStudy (amanda0618):

wheres the question?(laughing emoji)

OpenStudy (amanda0618):

I'm afraid to give you an answer cause i feel like it would be wrong...

OpenStudy (alekos):

I'd go with cc's 1 cubic centimetre = 1000 cubic millimetres

OpenStudy (anonymous):

i don't want the answer i want to know the operation to do

OpenStudy (phi):

they want you to write \(10^6 \) in "standard form" I 'm not sure what they mean by standard form, but they must have told you. It's probably just one million written as a number.

OpenStudy (phi):

can you figure out the length (which is also the width and height)?

OpenStudy (anonymous):

That i want to know

OpenStudy (phi):

the first part is probably just 1,000,000 You know the volume is 10*10 * 10*10 * 10*10 and you also know the volume is L*L*L (all 3 sides are the same length) so you should divide up 10*10 * 10*10 * 10*10 into 3 equal chunks. one chunk will be L

OpenStudy (anonymous):

so the answer is 333,333 cubic millimeters

OpenStudy (phi):

no, you divided 1000000/3 in other words, you found L so that L+L+L= 1,000,000 you want L*L*L= 1,000,000 notice that if you put parens in, like this: (10*10) * (10*10) * (10*10) = 1,000,000 that L matches up with 10*10 , which is 100

OpenStudy (anonymous):

Oh thanks phi for your help